OpenStudy (anonymous):

What is the exact diameter of an area rug with a circumference of thirty-six feet? eighteen pi feet eighteen divided by pi feet thirty-six divided by pi feet thirty-six pi feet @xanny6 @matricked @dg2

OpenStudy (anonymous):

given C=2*pi*r or C= pi*d (r is radius and d is diameter)

OpenStudy (anonymous):

so 36 =22d/7

OpenStudy (anonymous):

or 36 = d*pi so d= 36/pi

OpenStudy (anonymous):

So it's c?